At a local high school, GPA's are normally distributed with a mean of 2.5 and standard deviation of 0.6. What percentage of students at the high school have a GPA higher than 3.0?


1
Expert's answer
2021-10-17T18:08:46-0400

Let "X" be a random variable representing the the GPA's of students then, "X"~"N(\mu, \sigma^2)" where "\mu=2.5" and "\sigma^2=(0.6)^2." Therefore, "X"~"N(2.5,0.6^2)".

In order to determine the percentage of students who have a GPA higher than 3.0, we first find "p(X\gt3.0)".

Now,

"p(X\gt3)=p((X-\mu)/\sigma\gt(3-\mu)/\sigma)=p(Z\gt(3-2.5)/0.6)=p(Z\gt0.83)"

This can also be written as,

"p(Z\gt 0.83)=1-p(Z\lt0.83)=1-0.79673=0.20327"

The probability that students at the high school have a GPA higher than 3.0 is 0.20(2 decimal places).

Therefore, percentage of students who have a GPA higher than 3.0 is 0.20*100%=20%